# OBS-Diff: Accurate Pruning For Diffusion Models in One-Shot

### v1

@@ 1 Introduction
To improve the efficiency of the diffusion models, multiple research avenues have been proposed. A significant area of research is dedicated to expediting the sampling process by reducing the required denoising steps (Song et al., 2021; Lu et al., 2022) or knowledge distillation (Salimans & Ho, 2022; Sauer et al., 2024). Orthogonal to these efforts, model compression aims to reduce the intrinsic computational and memory footprint of the model itself. This category includes methods like quantization (He et al., 2023; Li et al., 2023b; Shang et al., 2023; Li et al., 2023a) and pruning, which is the primary focus of our work.
Pruning (Han et al., 2015; 2016; Wang et al., 2021; Fang et al., 2023a; Feng et al., 2024) is an effective way to reduce the computational and memory costs of deep neural networks. However, the rapid evolution of diffusion models underscores the severe limitations of existing pruning techniques. Current methods often lack generality (Fang et al., 2023b; Li et al., 2023c; Kim et al., 2024), as they are typically tailored to specific architectures like the U-Net and are not easily adapted to large-scale, text-to-image diffusion models with diverse structures (e.g., Multimodal Diffusion Transformer). Moreover, the efficiency gains from pruning are frequently undermined by computationally expensive requirements, such as the need for gradient information during pruning (Fang et al., 2023b; Zhang et al., 2024b) or a costly post-pruning fine-tuning stage. Furthermore, unstructured and semi-structured pruning remains largely unexplored for large-scale text-to-image diffusion models. All of these motivate our central research question: Can we develop a general and training-free pruning framework capable of pruning diffusion models with diverse architectures and supporting multiple pruning granularities in a one-shot manner?
@@ 2 Related Work
Several methods have explored pruning for diffusion models. An early approach, Diff-pruning (Fang et al., 2023b; a), introduced a gradient-based method for structured pruning. However, its demonstration on small-scale, non-text-to-image models (e.g., DDPMs) and its dependency on expensive retraining limit its applicability to modern, large-scale systems with diverse architectures. A significant line of research has since focused on compressing the UNet-based text-to-image diffusion models, with works like SnapFusion (Li et al., 2023c), MobileDiffusion (Zhao et al., 2024), BK-SDM (Kim et al., 2024), LAPTOP-Diff (Zhang et al., 2024a), and LD-Pruner (Castells et al., 2024) all targeting less salient components of the UNet architecture.
@@ 5.1 Settings
For text-to-image models, we compare against methods adapted from the Large Language Model (LLM) domain for unstructured/semi-structured sparsity, namely Wanda (Sun et al., 2024) and DSnoT (Zhang et al., 2024c), as well as standard magnitude pruning. For structured pruning, we employ an L1-norm based baseline (Li et al., 2017). On the CIFAR-10 DDPM, our method is directly compared with Diff-Pruning (Fang et al., 2023b). The sparsity refers to the pruning ratio of all the linear layers within MHA and FFN for each MMDiT block. For calibration, we utilize text prompts from the GCC3M dataset (Sharma et al., 2018). To ensure a fair comparison, all methods and baselines utilize identical configurations (computational resources provided in Appendix B).
@@ 5.4 Results On Structured Pruning
The results are presented in Table 3. The baseline $L_{1}$-norm pruning suffers from catastrophic performance degradation even at a modest 15% sparsity, with its FID score deteriorating from 31.59 to 158.89. In stark contrast, our method, OBS-Diff, demonstrates remarkable resilience. At the same 15% sparsity, OBS-Diff maintains an FID of 32.64, nearly identical to the dense model’s performance. This robustness persists up to 30% sparsity, where OBS-Diff sustains a strong FID of 34.51 while the baseline model fails completely (FID of 327.48). These findings highlight OBS-Diff’s superior ability to preserve critical model structures under aggressive structured pruning.
To benchmark our method against established techniques, we compare OBS-Diff with Diff-Pruning (Fang et al., 2023b), a well-recognized method that leverages gradient information for structured pruning on small class-conditional DDPMs. For brevity, the detailed results and analyses for the comparison are deferred to the  C.3, where our method outperforms consistently.
@@ C.2 Structured Pruning For SD3-medium
SD 3-medium Dense Model 2.03 B 2.84 31.59 0.3156 0.7549
@@ C.3 Comparison with Diff-Pruning on DDPM
Appendix D Additional Qualitative Results

### v2

@@ 1 Introduction
To improve the efficiency of the diffusion models, multiple research avenues have been proposed. One major line of work focuses on accelerating the sampling process by reducing the number of denoising steps (Song et al., 2021; Lu et al., 2022; Wang & Wang, 2025) or knowledge distillation (Salimans & Ho, 2022; Sauer et al., 2024). Orthogonal to these efforts, model compression aims to reduce the intrinsic computational and memory footprint of the model itself. This category includes methods like quantization (He et al., 2023; Li et al., 2023b; Shang et al., 2023; Li et al., 2023a) and pruning, which is the primary focus of our work.
The rapid evolution of diffusion models underscores the severe limitations of existing pruning techniques. Current methods often lack generality (Fang et al., 2023; Li et al., 2023c; Kim et al., 2024), as they are typically tailored to specific architectures like the U-Net and are not easily adapted to large-scale, text-to-image diffusion models with diverse structures (e.g., Multimodal Diffusion Transformer). Moreover, the efficiency gains from pruning are frequently undermined by computationally expensive requirements, such as the need for gradient information during pruning (Fang et al., 2023; Zhang et al., 2024b) or a costly post-pruning fine-tuning stage. Furthermore, unstructured and semi-structured pruning remains largely unexplored for large-scale text-to-image diffusion models. All of these motivate our central research question: Can we develop a general and training-free pruning framework capable of pruning diffusion models with diverse architectures and supporting multiple pruning granularities in a one-shot manner?
@@ 2 Related Work
Several methods have explored pruning for diffusion models. An early approach, Diff-pruning (Fang et al., 2023), introduced a gradient-based method for structured pruning. However, its demonstration on small-scale, non-text-to-image models (e.g., DDPMs) and its dependency on expensive retraining limit its applicability to modern, large-scale systems with diverse architectures. A significant line of research has since focused on compressing the UNet-based text-to-image diffusion models, with works like SnapFusion (Li et al., 2023c), MobileDiffusion (Zhao et al., 2024), BK-SDM (Kim et al., 2024), LAPTOP-Diff (Zhang et al., 2024a), and LD-Pruner (Castells et al., 2024) all targeting less salient components of the UNet architecture.
@@ 4 Methodology
We propose OBS-Diff, a one-shot, training-free pruning framework tailored for diffusion models. As illustrated in Figure 2, our method partitions the model into sequential “Module Packages” to amortize calibration costs. Within each package, we employ a novel Timestep-Aware Hessian construction to prioritize early denoising steps, enabling the simultaneous pruning of all target layers within the current package.
@@ 5.1 Settings
For text-to-image models, we compare against methods adapted from the Large Language Model (LLM) domain for unstructured/semi-structured sparsity, namely Wanda (Sun et al., 2024) and DSnoT (Zhang et al., 2024c), as well as standard magnitude pruning. For structured pruning, we employ an L1-norm based baseline (Li et al., 2017) and EcoDiff (Zhang et al., 2024b). On the CIFAR-10 DDPM, our method is directly compared with Diff-Pruning (Fang et al., 2023). The sparsity refers to the pruning ratio of all the linear layers within MHA and FFN for each MMDiT block. For calibration, we utilize text prompts from the GCC3M dataset (Sharma et al., 2018). To ensure a fair comparison, all methods and baselines utilize identical configurations (computational resources provided in Appendix B).
@@ 5.4 Results On Structured Pruning
Table 3: Performance of structured pruning on the SDXL (U-Net) model across various sparsity levels. Comparison includes the L1-norm baseline, EcoDiff, and our proposed OBS-Diff. The TFLOPs metric represents the theoretical computational cost for a single forward pass of the entire UNet. For each sparsity group, the best result per metric is highlighted in bold.
Model Sparsity Method #Params TFLOPs $\downarrow$ FID $\downarrow$ CLIP Score $\uparrow$ ImageReward $\uparrow$
SDXL Dense Model 2.57 B 5.98 29.21 0.3213 0.7635
15% $L_{1}$-norm 2.24 B 5.33 ($\downarrow$10.87%) 71.78 0.3035 -0.0006
EcoDiff 34.18 0.3100 -0.1870
OBS-Diff (Ours) 29.08 0.3215 0.6877
20% $L_{1}$-norm 2.13 B 5.12 ($\downarrow$14.38%) 133.07 0.2825 -0.7897
EcoDiff 42.98 0.2993 -0.6172
OBS-Diff (Ours) 29.19 0.3212 0.6461
30% $L_{1}$-norm 1.91 B 4.70 ($\downarrow$21.40%) 170.68 0.2711 -1.1694
EcoDiff 101.96 0.2465 -1.9161
OBS-Diff (Ours) 29.75 0.3204 0.4909
@@
The results are presented in Table 4 and Table 3. The baseline $L_{1}$-norm pruning suffers from catastrophic performance degradation even at a modest 15% sparsity, with its FID score deteriorating from 31.59 to 158.89 on SD 3.5-Large. In stark contrast, our method, OBS-Diff, demonstrates remarkable resilience. At the same 15% sparsity, OBS-Diff maintains an FID of 32.64, nearly identical to the dense model’s performance. This robustness persists up to 30% sparsity, where OBS-Diff sustains a strong FID of 34.51 while the baseline model fails completely (FID of 327.48). These findings highlight OBS-Diff’s superior ability to preserve critical model structures under aggressive structured pruning.
To benchmark our method against established techniques, we incorporate the comparison with EcoDiff (Zhang et al., 2024b), a state-of-the-art structured pruning framework for text-to-image diffusion models, directly into the main experiments. As shown in Table 3 and Table 4, while EcoDiff generally outperforms the naive $L_{1}$-norm baseline on SDXL, it still exhibits significant performance degradation compared to our method on both tables, especially at higher sparsity levels. For instance, on the U-Net based SDXL model (Table 3), EcoDiff yields an FID of 101.96 at 30% sparsity, whereas OBS-Diff achieves a substantially better FID of 29.75. This confirms that OBS-Diff generalizes effectively across diverse architectures, outperforming baselines on both MMDiT (SD 3.5) and U-Net (SDXL) backbones.
Finally, we compare OBS-Diff with Diff-Pruning (Fang et al., 2023), a well-recognized method that leverages gradient information for structured pruning on small class-conditional DDPMs. The detailed results for this specific comparison are deferred to the Appendix C.3, where our method outperforms Diff-Pruning consistently.
@@ C.2 Structured Pruning For SD3-medium
SD 3-medium Dense Model 2.03 B 2.84 36.14 0.3162 0.9029
@@ C.3 Comparison with Diff-Pruning on DDPM
Appendix D Robustness and Generalization Analysis
To demonstrate that our calibration (using only 100 prompts) does not overfit, we evaluated the fixed pruned model (SD3-Medium, 50% Unstructured) under inference conditions significantly different from the calibration settings (CFG 7.0, Steps 25, Euler). For fast evaluation, we evaluated all tasks on the MSCOCO 2014 validation 1K subset using the ImageReward metric.
D.1 Robustness to CFG Scales
We evaluated the pruned model across varying Classifier-Free Guidance (CFG) scales. As shown in Table 10, the pruned model achieves higher performance at CFG 9.0 (0.7044) than at the calibration setting of CFG 7.0 (0.6425). This indicates that our pruning strategy effectively preserves the model’s semantic generation capabilities and generalizes exceptionally well to higher guidance scales, which are critical for high-quality text-to-image synthesis.
Table 10: Robustness of the pruned SD3-Medium (50% Unstructured) across different CFG scales. The model was calibrated at CFG 7.0.
CFG Dense (ImageReward) Pruned (ImageReward) Performance
5.0 0.8319 0.5297 -
7.0 (Calibrated) 0.8510 0.6425 Baseline
9.0 0.8275 0.7044 Improved
D.2 Robustness to Sampling Steps
To address the concern regarding step counts, we evaluated the pruned model (calibrated at 25 steps) across 15, 25, and 50 inference steps. The results are summarized in Table 11. Although calibrated at 25 steps, the pruned model effectively leverages additional compute at 50 steps to generate higher-quality images. This confirms the pruning preserves the integrity of the underlying ODE trajectory.
Table 11: Robustness across varying inference sampling steps. The model was calibrated at 25 steps.
Steps Dense (ImageReward) Pruned (ImageReward) Trend
15 0.6988 0.4883 Fast Preview
25 (Calibrated) 0.8510 0.6425 Baseline
50 0.9391 0.7153 Improved Quality
D.3 Robustness across Samplers
We evaluated generalization across different solvers on both SD3-Medium (MMDiT) and SD v2.1 (U-Net).
  • SD3-Medium: Calibrated on Euler (1st-order), the model generalizes zero-shot to Heun (2nd-order), showing significant quality gains.
  • SD v2.1: We applied 40% unstructured pruning (calibrated on PNDM). As shown in Table 12, the relative performance ranking of the samplers is preserved between the Dense and Pruned models (e.g., DPM++ remains the highest performing), indicating the pruning is solver-agnostic.
Table 12: Generalization across different samplers for SD3-Medium and SD v2.1.
Model Sampler ImageReward (Dense) ImageReward (Pruned)
SD3-Medium Euler (Calibrated) 0.8510 0.6425
Heun (2nd Order) 0.9200 0.7249
SD v2.1 PNDM (Calibrated) 0.3432 0.1782
DPM++ 0.3889 0.2246
EDM 0.3442 0.1534
DDIM 0.3439 0.1579
D.4 Generalization to Out-of-Distribution (OOD) Prompts
We address the concern regarding calibration data in two ways:
  1. 1.
    Experimental Design: All main results in the paper already use GCC3M for calibration and MS-COCO for evaluation, representing a standard OOD setting.
  2. 2.
    New Validation Experiment: To rigorously test this, we calibrated two separate models—one using MS-COCO 2014 Train (In-Distribution) and one using GCC3M (Out-of-Distribution)—and evaluated both on the MS-COCO 2014 validation 5K subset.
As shown in Table 13, the performance is nearly identical. The model calibrated on OOD data (GCC3M) performs on par with (and even slightly better in FID than) the ID model. This definitively proves that OBS-Diff captures generalizable features and does not overfit to the calibration prompts.
Table 13: Comparison of models calibrated on In-Distribution (MS-COCO) vs. Out-of-Distribution (GCC3M) datasets, evaluated on MS-COCO validation set.
Calibration Dataset Evaluation FID $\downarrow$ CLIP Score $\uparrow$ ImageReward $\uparrow$
MS-COCO 2014 Train In-Distribution (ID) 27.93 0.3169 0.6547
GCC3M Train Out-of-Distribution (OOD) 27.20 0.3167 0.6468
Appendix E Additional Qualitative Results